# Focus on This, Not That! Steering LLMs With Adaptive Feature Specification

### v3

@@ 3 Methodology
We define $p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})$ by placing a small probability mass on the empty focus instruction prompt $\emptyset$ in order to aid in learning the underlying task, and then uniformly distribute the remaining probability mass over the remaining non-empty focus instructions. The objective in Equation 8 can be optimised through sampling using stochastic gradient descent (SGD) with popular optimisers such as AdamW (Loshchilov & Hutter, 2019). Further details on FT optimisation are provided in Appendix C.
Evaluating FIT under spurious correlations. After introducing FIT above, we now turn to settings where we can empirically train and evaluate it. A key aspect of our evaluation is the use of known spurious correlations, which simulate real-world scenarios where models can be misled by features that are spuriously predictive of the output label. By adjusting the co-occurrence rate between spurious features and their associated labels, we can test FIT’s ability to dynamically steer a model’s responses depending on the features on which it is focusing or ignoring.
We define the co-ocurrence rate, or predictivity (Hermann et al., 2024), between spurious feature values and the label with which they are spuriously correlated by $\rho_{\text{spurious}}$. Specifically:
Definition 3.1.
@@
(Defining $\rho_{\text{spurious}}$) . Let $S\in\mathcal{S}\subseteq\mathcal{F}$ denote a spurious feature. Suppose that a value of $S$, say $s\in\text{Val}(S)$, is spuriously correlated with label $y_{s}$. Then, for ground truth dataset label $Y$, we define
$\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$ (9)
@@
By varying $\rho_{\text{spurious}}(s)$, we can control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well as core features under distribution shift.
To examine how well models can utilise focus instructions for steerable behaviour in a controlled environment, we construct synthetic datasets such that the spurious feature $S$ is not predictive of the ground-truth label $Y$ (i.e., $Y\perp\!\!\!\perp S$) and the core feature $C$ is not predictive of the spurious label $Y_{S}$ (i.e., $Y_{S}\perp\!\!\!\perp C$). We enforce these conditions by setting $\rho_{\text{spurious}}=1/N$ (for $N$ label classes) and ensuring a balanced label distribution during training. In Appendix I and Appendix K, we verify that our synthetic SS and SMNLI training sets introduced in Section 4.1 and Figure 4 respectively, satisfy these independence assumptions.
@@
Next, we evaluate FIT across several test sets with varying predictivity levels:
@@
We further evaluate FIT under another form distribution shift, specifically on our SMNLI dataset (c.f. Section 4.2).. Here, the specific values taken by spurious features do not overlap between the training and test sets.
  • $\mathcal{D}^{s}_{\cdot}$: Test datasets where the spurious feature values are distinct from those within the training set. Here, we use the same predictivity levels as in the initial datasets presented above.
@@ 4 Experiments
Before reporting the main results, we introduce the evaluation metric (focus accuracy) that we report, baselines, models, and training settings used throughout the experiments. In Section 4.1, we first verify that FIT performs well on the simpler SS dataset, a synthetic sentiment analysis dataset derived from SST-5 (Socher et al., 2013b). We then demonstrate in Section 4.2 that FIT generalises to more complex features and handles feature-value distribution shifts on the SMNLI dataset, a sub-sampled version of the MNLI dataset (Williams et al., 2018). Finally, in Section 4.3, we show that FIT has practical, real-world impact by effectively mitigating bias in the BBQ dataset (Parrish et al., 2022), where we further illustrate FIT’s ability to generalise to new features seen for the first time when performing inference.
Although the primary focus of FIT is on adaptively steering LLMs at inference time, which is what we focus on in this paper, we include an additional debiasing experiment for comparison on the BBQ dataset in Appendix H for completeness. While bias mitigation is a valuable and natural application of FIT, it is not its primary objective. Instead, this inclusion highlights FIT’s broader utility as a tool for model control and adaptability, demonstrating that it performs on par with dedicated bias mitigation techniques while also offering the unique advantage of test-time steerability.
Metrics. We define the focus accuracy for a focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ as the proportion of samples where the model’s prediction aligns with the focus label, $y_{\text{focus}}$, as specified in Equation 3. Specifically, for each sample $(x,y)\in\mathcal{D}$, the model produces a prediction $\hat{y}\sim p_{\theta}(\cdot\mid I,I_{\text{focus}},x)$ based on a fixed focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. The focus label, $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},s,y)$, corresponds to the target output given the focus instruction for the input $x$ with ground truth label $y$, a spurious feature value $s$ present in $x$. Focus accuracy for focus instruction $I_{\text{focus}}$, denoted $\mathcal{A}_{\text{focus}}(I_{\text{focus}})$, is computed as the fraction of correct predictions with respect to the focus label:
@@
We report focus accuracy for each model on all dataset splits, using the prompt types and focus instructions detailed in Appendix E. Generations are evaluated through simple pattern matching due to the use of constrained beam decoding (Anderson et al., 2017). Further details are provided in Appendix D.
@@
Models and training settings. We evaluate FIT using three popular LLMs that span a range of model sizes: Llama-3.1-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.3  (Jiang et al., 2023), and Vicuna-13B-v1.5  (Chiang et al., 2023). The models are fine-tuned using parameter-efficient SFT with LoRA (Hu et al., 2021), leveraging Hugging Face’s SFTTrainer (Wolf et al., 2020) with default hyperparameters. Early stopping is applied based on validation loss, as defined in Equation 8. For generation, we use constrained beam decoding (Anderson et al., 2017) and use fully verbalised (natural language) labels during both training and testing, except for the multi-choice BBQ dataset. For further training details, refer to Appendix C.
Figure 3: SS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the SS dataset.
@@
Baselines. We compare against the following baselines in the main section of the paper: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline, $\text{SFT}(y_{\text{focus}})$, follows the same setup as the FIT method (trained on sampled inputs and focus labels), but without the inclusion of focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions in FIT. The few-shot baseline involves using 5 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. In Appendix F, we include two additional baselines: zero-shot and vanilla SFT for a more complete comparison with FIT. Further details of baselines and their results in comparison to FIT can be found in Appendix F.
4.1 Validation of FIT on the SS dataset
@@
Figure 4: SMNLI focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT models on the (a) SMNLI standard test sets $\mathcal{D}$, and (b) SMNLI test sets under feature value shift $\mathcal{D}^{s}$.
@@
Results. Figure 3 (a) shows the focus accuracy results of the three LLMs on the SS dataset using few-shot prompting, after $\text{SFT}(y_{\text{focus}})$ and after FIT. We see that across all focus instructions and all models, FIT shows significant improvement over the baselines, achieving very high focus accuracy across all focus instruction types and across all test sets with varying predictivity levels.
Key takeaways. High focus accuracy on SS indicates that FIT successfully steers model responses based on the feature on which it is instructed to focus or to not focus on.
4.2 FIT performs well with more complex features on the SMNLI dataset and generalises under distribution shift
[…]
Moreover, for SMNLI, we hold out specific genres at test time to evaluate our model’s ability to generalise under distribution shift when feature values change. We do this by sub-sampling a held-out portion of the MNLI dataset. During training, we use three selected genres (government, fiction, and travel) to evaluate our models. At test time, we additionally add three held-out genres (facetoface, nineeleven, and verbatim). We again ensure that $\rho_{\text{spurious}}$ is constant within each dataset split for all feature values, and use the same set of corresponding $\rho_{\text{spurious}}$ as within the SMNLI test sets described above. Further details of the SMNLI dataset can be found in Appendix K.
Results. Figure 4 (a) depicts the focus accuracy results of the three models on the SMNLI test splits. We observe that for the more complex feature of genre, FIT achieves very high focus accuracy, significantly improving over the baselines. This demonstrates that FIT effectively trains the model to handle more complex features, allowing it to dynamically focus on or disregard these features when making predictions.
Figure 4 (b) shows the focus accuracy of models on the feature-shifted test sets. When focusing on the core feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the third row of Figure 4) (b). Note that, while we observe low focus accuracy when focusing on spurious features, this is expected, as the spurious labels associated with these new genres were not encountered during training. Thus when focusing on these features the model does not know what label to predict. This result highlights that the focus-tuned models have indeed learned spurious associations during training and correctly reproduces them when instructed to focus on these spurious features, even for new spurious feature values. When instructed to focus on the core feature or when instructed to ignore the spurious feature, the model still shows strong generalisation in the presence of distribution shift.
Key takeaways. FIT achieves high focus accuracy on more complex features and maintains strong performance under distribution shift in terms of feature values. This demonstrates FIT’s ability to generalise to new contexts and reliably handle changing feature values.
[…]
Figure 5: BBQ focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the BBQ dataset.
@@ 4.3 FIT steers behaviour in the presence of social bias data and generalises to unseen features
Key takeaways. FIT can effectively teach models to adjust their responses based on knowledge of social biases. This ability generalises to biases not seen during FIT training, indicating FIT’s utility for bias mitigation.
5 Ablation
Generalisation to different test-time prompt formats. As observed in the IT literature, instruction-tuned models sometimes memorise instruction formats and struggle to follow paraphrased instructions at test time (Ghosh et al., 2024). In Appendix G (Figure 9), we compare the performance of models on the SMNLI dataset when using the same focus instructions at training and test time versus using paraphrased instructions at test time. We generate 10 different test-time focus instructions of each instruction type defined in Equation 2 by paraphrasing the existing focus instruction using ChatGPT (OpenAI, 2022). The results show minimal variation in focus accuracy across different dataset splits and focus features, even when testing on paraphrased prompts, indicating that FIT indeed teaches models a general capacity to focus on or ignore features regardless of the specific way that focus instructions are phrased.
Instruction Following After FIT. Prior studies suggest that SFT can impair the instruction-following abilities of LLMs (Fu et al., 2024; Dou et al., 2024). We evaluate whether FIT impacts instruction adherence by comparing pre-trained and FIT models trained on the SMNLI dataset (see Section 4.2). Using 500 samples from the Alpaca-GPT instruction-tuning dataset (Peng et al., 2023), responses were rated by GPT-4o (Achiam et al., 2023) on a 1–5 scale, where 5 denotes perfect alignment and 1 indicates a completely incorrect response.
LLaMA Mistral Vicuna
Pre-Trained Avg. Rating 3.51 3.65 3.46
FIT Avg. Rating 3.45 3.65 3.50

### v4

@@ 3.2 Focus Instruction Tuning (FIT)
We define $p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})$ by placing a small probability mass on the empty focus instruction prompt $\emptyset$ in order to aid in learning the underlying task, and then uniformly distribute the remaining probability mass over the remaining non-empty focus instructions. The objective in Equation 3 can be optimised through sampling using stochastic gradient descent (SGD) with popular optimisers such as AdamW (Loshchilov & Hutter, 2019). Further details on FT optimisation are provided in Appendix D.
3.3 Evaluating FIT Under Controlled Spurious Correlations on Synthetic Datasets
Before turning to real-world data (see Section 4.2), we first train and evaluate FIT in a fully controlled setting. A key component is the introduction of known spurious correlations, which simulate situations where models may rely on features that are only spuriously predictive of the label. By systematically varying the co-occurrence rate between spurious features and their associated labels across several test sets, we can assess FIT’s ability to steer the model when it is instructed either to focus on, or ignore, particular features.
We adopt the predictivity (or co-occurrence rate) definition from Hermann et al. (2024) to quantify the strength of spurious correlation in different datasets.
Definition 2 (Predictivity, $\rho_{\text{spurious}}$).
@@ 3.3 Evaluating FIT Under Controlled Spurious Correlations on Synthetic Datasets
Let $S\in\mathcal{S}\subseteq\mathcal{F}$ be a spurious feature. Assume that a concrete value $s\in\operatorname{Val}(S)$ is spuriously correlated with label $y_{s}\in\mathcal{Y}$. We define its predictivity
$\rho_{\text{spurious}}(s)\;=\;\mathbb{P}(Y=y_{s}\mid S=s\bigr{)},$ (4)
where $Y$ is the ground-truth label random variable.
@@
By varying $\rho_{\text{spurious}}(s)$, we can precisely control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well as core features under distribution shift.
Synthetic Training Conditions.
During training we construct datasets so that spurious features $S$ are independent of the ground-truth label $Y$ ($Y\perp\!\!\!\perp S$) and, symmetrically, that the core feature $C$ is independent of the spurious-label variables $Y_{S}$ ($Y_{S}\perp\!\!\!\perp C$). We enforce this by setting $\rho_{\text{spurious}}(s)=1/N$ for every $s\in\operatorname{Val}(S)$, where $N=|\mathcal{Y}|$ is the number of classes for a given task. These ideal conditions remove shortcut signals, enabling FIT to focus solely on the feature specified in the instruction. However, Section 4.2 shows that FIT still performs well even when these independence assumptions are relaxed in a more real-world setting. See Appendix D for a more detailed discussion on the independence assumptions above.
In Appendix F and Appendix E, we empirically verify that the training splits of both our synthetic SMNLI dataset (introduced in Section 4.1) and the additional sentiment-analysis dataset SS indeed satisfy the independence constraints described above.
@@
Synthetic Test Sets. We evaluate FIT across several test sets with varying predictivity levels:
@@
We further evaluate FIT under another form of distribution shift specifically on our SMNLI dataset (c.f. Section 4.1). Here, the specific values taken by spurious features do not overlap between the training and test sets.
  • $\mathcal{D}^{s}_{\cdot}$: Test datasets where the spurious feature values are distinct from those within the training set and the test sets above. Here, we use the same predictivity levels as in the initial datasets presented above.
@@ 4 Experiments
Before reporting the main results, we introduce the evaluation metric (focus accuracy) that we report, baselines, models, and training settings used throughout the experiments. We first demonstrate in Section 4.1 that FIT generalises to subtle textual features and handles feature-value distribution shifts on the SMNLI dataset, a sub-sampled version of the MNLI dataset (Williams et al., 2018). In Appendix E, we additionally verify that FIT performs well on the SS dataset, a synthetic sentiment analysis dataset derived from SST-5 (Socher et al., 2013b). Finally, in Section 4.2, we show that FIT has practical, real-world impact by effectively mitigating bias in the BBQ dataset (Parrish et al., 2022), where we further illustrate FIT’s ability to generalise to new features seen for the first time when performing inference.
Although the primary focus of FIT is on adaptively steering LLMs at inference time, which is what we focus on in this paper, we include an additional debiasing experiment for comparison on the BBQ dataset in Appendix I for completeness. While bias mitigation is a valuable and natural application of FIT, it is not its primary objective. Instead, this inclusion highlights FIT’s broader utility as a tool for model control and adaptability, demonstrating that it performs on par with dedicated bias mitigation techniques while also offering the unique advantage of test-time steerability.
Metrics. We define the focus accuracy for a focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ as the proportion of samples where the model’s prediction aligns with the focus label, $y_{\text{focus}}$, as specified in Definition 1.
Definition 3 (Focus Accuracy, $\mathcal{A}_{\text{focus}}$).
For a sample $(x,y)\sim p_{\text{data}}$ and a fixed focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$, we consider a model’s prediction of the focus label $\hat{y}\sim p_{\theta}(\cdot|I,I_{\text{focus}},x)$. Focus accuracy for focus instruction $I_{\text{focus}}$, denoted $\mathcal{A}_{\text{focus}}(I_{\text{focus}})$, is defined as
@@
We report focus accuracy for each model on all dataset splits, using the prompt types and focus instructions detailed in Appendix C. Generations are evaluated through simple pattern matching due to the use of constrained beam decoding (Anderson et al., 2017). See Section D.2 for further details.
@@
Models and Training Settings. We evaluate FIT using three popular LLMs that span a range of model sizes: Llama-3.1-8B-Instruct  (Dubey et al., 2024), Mistral-7B-Instruct-v0.3  (Jiang et al., 2023), and Vicuna-13B-v1.5  (Chiang et al., 2023). The models are fine-tuned using parameter-efficient SFT with LoRA (Hu et al., 2021), leveraging Hugging Face’s SFTTrainer (Wolf et al., 2020). Early stopping is applied based on validation loss, as defined in Equation 3. For generation, we use constrained beam decoding (Anderson et al., 2017) and use fully verbalised (natural language) labels during both training and testing, except for the multi-choice BBQ dataset. Focus accuracies are reported over four independent repeats for each experiment. For further training details, refer to Appendix D.
@@
Baselines. We compare against the following baselines in the main section of the paper: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline, $\text{SFT}(y_{\text{focus}})$, follows the same setup as the FIT method (trained on sampled inputs and focus labels), but without the inclusion of focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions.
Recent findings (Wu et al., 2025) indicate that LSMs significantly underperform compared to SFT methods. Hence, we include SFT baselines rather than LSM baselines within this work. The few-shot baseline involves using 4 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. In Section D.4, we detail and include two additional baselines: zero-shot and vanilla SFT for a more complete comparison with FIT.
@@
Figure 3: SMNLI Focus Accuracies ($\uparrow$). Mean focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT models on the (a) SMNLI standard test sets $\mathcal{D}$, and (b) SMNLI test sets under feature value shift $\mathcal{D}^{s}$. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$ and few-shot methods across models and $\mathcal{I}_{\text{focus}}$ are 6.47%, 7.98% and 0.500% respectively. fcs = focus, ign = ignore.
4.1 FIT Performs Well on the SMNLI Dataset and Generalises Under Distribution Shift
We evaluate our method on a dataset with subtle textual features. Specifically, we construct an NLI dataset by sub-sampling from MNLI (Williams et al., 2018), where we induce a spurious correlation between text genres and labels through our subsampling process. We call this subsampled dataset the SMNLI dataset.
Dataset Construction. Figure 4 illustrates the data generating process (DGP) describing how we subsample examples to induce spurious correlations between feature value $s\in\text{Val}(S)$ and a particular associated label $y_{s}\in\text{Val}(Y)$. The feature set that we consider is defined as =$\mathcal{F}=\{C,S\}$, where $C$ is the NLI relationship and $S$ is the genre of a given premise-hypothesis pair.
The co-occurrence rate of genres and their spuriously associated labels is governed by $\rho_{\text{spurious}}$, which varies across the test sets discussed in Section 3. We ensure that $\rho_{\text{spurious}}$ is the same for all feature values in $\text{Val}(S)$ within each dataset split. In particular, we set $\rho_{\text{spurious}}$ to be $1/3,1/3,0.9,0.1$ and $0.9$ on $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{iid}}$, $\mathcal{D}_{\text{high}}$, $\mathcal{D}_{\text{low}}$ and $\mathcal{D}_{\text{flipped}}$ respectively.
Figure 4: SMNLI DGP. DGP describing the subsampling process of MNLI to introduce the spurious genre feature $S$. Here, $\mathcal{U}(\mathcal{S})$ is the uniform distribution over genres, $\mathrm{Val}(C)=\{0,1,2\}$ are the NLI labels (with $y_{S}$ tied to each $S$), and $p_{\mathcal{D}}(\cdot|C,S)$ is the MNLI conditional distribution over premise–hypothesis pairs.
Moreover, we hold out specific genres at test time to evaluate our model’s ability to generalise under distribution shift when feature values change. We do this by sub-sampling a held-out portion of the MNLI dataset. During training, we use three selected genres $\mathcal{S}=\{\text{government},\;\text{fiction},\;\text{travel}\}$ to train and evaluate our models. We additionally add three held-out genres $\mathcal{S}=\{\text{facetoface},\;\text{nineeleven},\;\text{verbatim}\}$. We again ensure that $\rho_{\text{spurious}}$ is constant within each of these shifted splits across feature values, and use the same set of corresponding $\rho_{\text{spurious}}$ as within the SMNLI test sets described above. This process is again governed by a DGP shown in Figure 4, giving us precise control over the data synthesis process for the SMNLI dataset. Further details of the SMNLI dataset can be found in Appendix F.
Results. Figure 3 (a) depicts the focus accuracy results of the three models on the SMNLI test splits. We observe that for both the core and genre feature, FIT achieves very high focus accuracy, significantly improving over the baselines. This demonstrates that FIT effectively trains the model to handle subtle textual features, allowing it to dynamically focus on or disregard these features when making predictions.
Figure 3 (b) shows the focus accuracy of models on the feature-shifted test sets. When focusing on the core feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the third row of Figure 3 (b)), generalising generally more robustly than the baseline methods when focus is over core features.
While we observe low focus accuracy when focusing on spurious features, this is expected because the spurious labels associated with these new genres were not encountered during training, so the model cannot know these new relations. Importantly, FIT remains steerable, changing its predictions depending on what is focused on, and this holds steady across all predictivity levels for the new spurious genres. In contrast, the baselines show decreasing focus accuracy as predictivity decreases, indicating a tendency to predict the causal label under distribution shift. This shows that these models do not change their behaviour when instructed to change their focus, and thus have poorer steering ability under distribution shift compared to FIT.
Key Takeaways. FIT achieves strong steerability, which is maintained under distribution shift. This demonstrates FIT’s generalisation to new contexts with changing feature values.
Bias Benchmark for QA (BBQ) Dataset. We experiment with BBQ (Parrish et al., 2022), a MCQA benchmark annotated with social biases that are relevant to any given answer, such as stereotypes that would imply a given answer to an otherwise ambiguous question (see Figure 1).
@@ 4.2 FIT Steers Behaviour in the Presence of Social Bias Data and Generalises to Unseen Features
Figure 5: BBQ Focus Accuracies ($\uparrow$). Mean focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the BBQ dataset. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$ and few-shot methods across models and $\mathcal{I}_{\text{focus}}$ are 4.07%, 10.7% and 0.600% respectively. fcs = focus, ign = ignore.
@@
Key Takeaways. FIT effectively teaches models to adjust their responses based on knowledge of social biases. This generalises to biases not seen during training, indicating FIT’s utility for bias mitigation.
5 Ablation Studies
5.1 Extending FIT to NLG Tasks.
The primary focus of our experiments has been on classification and MCQA datasets, due to the cost and difficulty of collecting high-quality natural language generation (NLG) benchmarks. As an initial step towards extending FIT to NLG tasks, we introduce BBQ-NLG, a dataset that follows the BBQ MCQA setup (see Figure 5) except where we now drop the fixed answer options for examples and require the model both to identify all plausible answers from context and to generate the correct answer in a fully verbalised form. To assess generation accuracy against ground-truth responses, we use a pre-trained Llama-3.1-8B-Instruct model as an automated judge. Further details and full results are given in Appendix H
The results shown in Figure 6 confirm that FIT can steer models effectively at inference time and generalise to novel, unseen features even in this NLG-style setting, underscoring that extending FIT to NLG tasks is a particularly promising avenue for future research.
Figure 6: BBQ-NLG FIT Focus Accuracies ($\uparrow$). Mean focus accuracy ($\mathcal{A}_{\text{focus}}$) of FIT models on the BBQ-NLG dataset. The maximum standard deviation across across FIT models and $\mathcal{I}_{\text{focus}}$ is . fcs = focus, ign = ignore.
5.2 Robustness to Prompt Phrasing.
Instruction-tuned models can overfit to the exact wording of their prompts, faltering on paraphrases (Ghosh et al., 2024). Figure 7 contrasts SMNLI focus accuracy with test-time focus instructions $\mathcal{I}_{\text{focus}}$ are either (i) the original training prompts given in Figure 9 or (ii) one of ten ChatGPT-paraphrased variants for each instruction type in Equation 2. Across splits and focus features, focus accuracy varies negligibly, indicating that FIT is robust to the particular phrasing of focus instructions.
Figure 7: Different Training and Test $\mathcal{I}_{\text{focus}}$ Focus Accuracy ($\uparrow$). SMNLI focus accuracies ($\mathcal{A}_{\text{focus}}$) when test focus instructions $I_{\text{focus}}$ prompts are drawn from the training focus instruction set (top) (see Figure 9) versus a paraphrased focus instruction set (bottom). fcs = focus, ign = ignore.
5.3 FIT does not Affect General Capabilities.
Prior work has shown that SFT can erode the instruction-following capabilities of LLMs (Fu et al., 2024; Dou et al., 2024). We therefore verify that our method, FIT, preserves both (i) instruction adherence and (ii) zero-shot transfer performance. All FIT models in this section are trained only on the SMNLI dataset (see Section 4.1).
Instruction Following (Alpaca-GPT Dataset).
We sample 500 prompts from the Alpaca-GPT set (Peng et al., 2023) and score each model’s response with GPT-4o (Achiam et al., 2023) on a 1–5 scale (5 = perfect alignment). Table 1 reports the mean score before and after FIT along with two-sided Wilcoxon signed-rank $p$-values. Across Llama, Mistral and Vicuna, the differences in ratings are small ($\leq 0.05$ points) and never significant ($p>0.05$), confirming that FIT preserves instruction-following capabilities.
Model Llama Mistral Vicuna
Pre-Trained Avg. Rating ($\uparrow$) 3.51 3.65 3.46
FIT Avg. Rating ($\uparrow$) 3.45 3.65 3.50